Carter Company sells a product for $100 per unit. The variable cost is $40 per unit, while fixed costs are $300,000. Additionally, the income tax rate is 40 percent.

Required:

  1. Calculate the contribution margin ratio (round your answer to xx.x %).
  2. Calculate the break-even point in sales units.
  3. Calculate the break-even point in sales dollars or revenues.
  4. How many units need to be sold to generate a pretax income of $180,000?
  5. Recalculate the break-even point in sales units if the selling price increased to $120 per unit. Round your answer to the nearest whole number.
  6. Calculate the after-tax net income assuming that 6,000 units are sold.
Answer #1

a.Contribution Margin Ratio = (Selling price – variable cost)/Selling price

= (100-40)/100

= 60%

b.Break even point in units = Total fixed costs/Contribution Margin per unit

= 300,000/60

= 5,000 units

c.Break even point in sales dollars = Fixed costs/Contribution margin ratio

= 300,000/60%

= $500,000

d.Units required = (Desired Income + Fixed costs)/Contribution Margin per unit

= (180,000+300,000)/60

= 8,000 units

e.Break even units = 300,000/(120-40)

= 3,750 units

f.After tax income = (100-40)*6000 – 300,000

= $60,000